OpenStudy (anonymous):

Find an expression for the function whose graph is on the given curve. The bottom half of the parabola x+(y-3)^2=0

OpenStudy (anonymous):

The bottom half of the parabola is going to occur when we subtract whatever square root, instead of adding it, so we have \[\begin{align} &x+(y-3)^2=0\\ &(y-3)^2=-x\\ &y-3=-\sqrt{-x}\\ &y=3-\sqrt{-x} \end{align}\] If you're doing this for someone fussy, you should probably restrict the domain to \((-\infty,0]\).

OpenStudy (anonymous):

Thanks so much

OpenStudy (anonymous):

My pleasure.